OpenStudy (anonymous):

Solve the system using an equivalent system. -2x - y = 0 3x - y = 10

OpenStudy (anonymous):

Subtract one from the other to get rid of the y

OpenStudy (anonymous):

subtract -2 and 3?

OpenStudy (anonymous):

I mean subtract one equation (the whole thing) from the other one

OpenStudy (anonymous):

ok so i did -2 - 0 which is -2 and i did 3 - 10 which is -7

OpenStudy (anonymous):

did i do it right? these are my options (4, –2) (–2, 4) (2, –4) (–4, 2)

OpenStudy (anonymous):

-2x - y = 0 minus 3x - y = 10 equals ----------- So -2x - 3x -y -(-y) = 0-10

OpenStudy (anonymous):

i got -5

OpenStudy (anonymous):

-2x-3x = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Nyet! -2x -3x = -5x What about -y -(-y) ?

OpenStudy (anonymous):

what are we putting as y?

OpenStudy (anonymous):

We are not putting anything as anything at the moment. All we are doing is subtracting one equation from the other one.

OpenStudy (anonymous):

We are trying to get this -2x - 3x -y -(-y) = 0-10 So far we have got -2x -3x = -5x and the next bit is -y - (-y) = ?

OpenStudy (anonymous):

-5 or y?

OpenStudy (anonymous):

-y - (-y) = -y +y = 0

OpenStudy (anonymous):

So -2x - 3x -y -(-y) = 0-10 -> -5x = -10 so what is x?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

-5x = -10 Put x = 4

OpenStudy (anonymous):

ok so my answer is -2, and 4?